What is the velocity(speed) of a neutron (mass=1.67 ×10-24 g) whose wavelength is 0.155 nm?
Answer #1

de Broglie wavelength,

or,

where,

= wavelength associated with the particle = 0.155 nm = 0.155 x 10-9 m

h = Planck's constant = 6.62 x 10-34 J s

m = mass of the particle = 1.67 ×10-24 g = 1.67 ×10-27 kg

v = velocity of the particle

or,

    = 2557 m/s

Therefore, the velocity of the neutron = 2557 m/s